# Stacked Pseudo-Convergent Sequences and Polynomial Dedekind Domains

### v2

@@ Introduction
Here, given a prime $p\in\mathbb{Z}$ and the DVR $\mathbb{Z}_{(p)}$ of $\mathbb{Q}$, we are interested in describing residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, that is, valuation domains $W$ of $\mathbb{Q}(X)$ lying above $\mathbb{Z}_{(p)}$ such that the residue field extension $W/M\supseteq\mathbb{Z}/p\mathbb{Z}$ is algebraic and the value group $\Gamma_{w}$ of the associated valuation $w$ to $W$ is contained in the divisible hull of the value group of $\mathbb{Z}_{(p)}$ (i.e., the rationals). These valuation domains arise naturally as overrings of rings of integer-valued polynomials and Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ (see [9, 19]) and also in the description of closed subfields of $\mathbb{C}_{p}$ ([12]), the completion of an algebraic closure $\overline{\mathbb{Q}_{p}}$ of the field of $p$-adic numbers $\mathbb{Q}_{p}$. In the case $W$ is a DVR and the residue field extension is finite, then by [21, Theorem 2.5 & Proposition 2.2] there exists an element $\alpha$ in $\overline{\mathbb{Q}_{p}}$, transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\overline{\mathbb{Z}_{p}}\}$, where $\overline{\mathbb{Z}_{p}}$ is the absolute integral closure of $\mathbb{Z}_{p}$ (i.e., the integral closure of $\mathbb{Z}_{p}$ in $\overline{\mathbb{Q}_{p}}$; note that $\overline{\mathbb{Z}_{p}}$ is the valuation domain of the unique extension of $v_{p}$ to $\overline{\mathbb{Q}_{p}}$). In general, given a residually algebraic torsion extension $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, there exists a pseudo-convergent sequence $E=\{s_{n}\}_{n\in\mathbb{N}}$ in $\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{(p),E}=\{\phi\in\mathbb{Q}(X)\mid\phi(s_{n})\in\overline{\mathbb{Z}_{p}},\text{ for all sufficiently large }n\in\mathbb{N}\}$ (Proposition 2.24). One of the main result of this paper is to show that we may assume that $E$ is stacked (in a sense we make clear in §2; see Theorem 2.5). In particular, if $W$ is a DVR of $\mathbb{Q}(X)$ extending $\mathbb{Z}_{(p)}$ such that the extension of the residue fields is infinite algebraic, then there exists $\alpha$ in $\mathbb{C}_{p}\setminus\overline{\mathbb{Q}_{p}}$, such that $W=\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$, where $\mathbb{O}_{p}$ is the completion of $\overline{\mathbb{Z}_{p}}$ (equivalently, $\mathbb{O}_{p}$ is the valuation domain of the unique extension of $v_{p}$ to $\mathbb{C}_{p}$; see Corollary 2.28). Necessarily, the (transcendental) extension $\mathbb{Q}_{p}(\alpha)/\mathbb{Q}_{p}$ has finite ramification.
@@
It is worth recalling that in [4, §5, 1., & Theorem 5.1] a residually algebraic torsion extension $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ is realized as the limit of a sequence of residually transcendental extensions $W_{n}$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ (i.e., the residue field extension of $W_{n}$ over $\mathbb{Z}_{(p)}$ is transcendental); moreover, for each $n\in\mathbb{N}$, $W_{n}$ is defined by a minimal pair $(s_{n},\delta_{n})$ (as explained in [4, p. 282]; for the definition of minimal pair see §1.2). Here, $W$ is realized as the valuation domain $\mathbb{Z}_{(p),E}$, where for each $n\in\mathbb{N}$, $(s_{n},\delta_{n}=v_{p}(s_{n+1}-s_{n}))$ is a minimal pair, too.
@@
As an application of the above results, we are able to complete the classification of the family of Dedekind domains $R$ between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ started in [19]. In that paper we described the Dedekind domains of this family whose residue fields of prime characteristic are finite fields ([19, Theorem 2.17]); the description is obtained by means of the notion of rings of integer-valued polynomials over algebras. We also showed that, given a group $G$ which is the direct sum of a countable family of finitely generated abelian groups, there exists a Dedekind domain $R$ with finite residue fields of prime characteristic, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, with class group $G$ ([19, Theorem 3.1]).
@@ 1.2 Distinguished pairs
By Krasner’s Lemma, for each of the relevant $c$’s we have $v(a-c)\leq\omega_{K}(a)$. Note that the ramification index of $K(a,c)$ over $K$ is (strictly) bounded by $[K(a):K]^{2}$. In particular, since the value $v(a-c)$ belongs to $\Gamma_{a-c}$, it follows that there exists $N\in\mathbb{N}$, independent from each of the above $c$’s, such that $Nv(a-c)\in\Gamma_{v}\cong\mathbb{Z}$. Hence, the set $\{v(a-c)\mid c\in\overline{K},[K(c):K]<[K(a):K]\}$ (which is a subset of $\Gamma_{\overline{v}}$) is bounded from above and its elements have bounded torsion. It follows that this set has a maximum, which is equal to $\delta_{K}(a)$ by its very definition. ∎
@@
[…]
In other words, $(a,\delta)$ is a minimal pair if for every $b\in B(a,\delta)=\{x\in\overline{K}\mid v(a-x)\geq\delta\}$, we have $[K(b):K]\geq[K(a):K]$ (i.e., $a$ is a ’center’ of the ball $B(a,\delta)$ of minimal degree). By Lemma 1.2, $(a,\delta)$ is a minimal pair if and only if $\delta>\delta_{K}(a)$. In particular, if $\delta>\omega_{K}(a)$, then $(a,\delta)$ is a minimal pair.
@@
[…]
@@
[…]
@@
In particular, by the above Theorem, $\Gamma_{b}\subseteq\Gamma_{a}$, $k_{b}\subseteq k_{a}$ and $[K(b):K]\mid[K(a):K]$.
@@ 2 Stacked pseudo-convergent sequences of $\overline{\mathbb{Q}_{p}}$
Let $\mathbb{P}\subset\mathbb{Z}$ be the set of prime numbers and let $p\in\mathbb{P}$ be a fixed prime. We let $\mathbb{Z}_{(p)}$ be the localization of $\mathbb{Z}$ at the prime ideal $p\mathbb{Z}$, $\mathbb{Z}_{p}$ the ring of $p$-adic integers and $\mathbb{Q}_{p}$ its field of fractions, the field of $p$-adic numbers. If $v_{p}$ denotes the usual $p$-adic valuation, then $\mathbb{Z}_{p}$ ($\mathbb{Q}_{p}$, respectively) is the completion of $\mathbb{Z}$ ($\mathbb{Q}$, respectively) with respect to the $p$-adic valuation. We denote by $\overline{\mathbb{Q}_{p}}$ a fixed algebraic closure of $\mathbb{Q}_{p}$ and still denote the unique extension of $v_{p}$ to $\overline{\mathbb{Q}_{p}}$ by $v_{p}$. Note that $\overline{\mathbb{Q}_{p}}$ is a rank one non-discrete valued field with valuation domain denoted by $\overline{\mathbb{Z}_{p}}$, the integral closure of $\mathbb{Z}_{p}$ in $\overline{\mathbb{Q}_{p}}$. We will use the well-known fact that $\mathbb{Q}_{p}$ has only finitely many extensions of a given degree (see for example [17, Corollary 2, Chapter V, p. 202]).
@@ 2.1 Residually algebraic torsion extensions of $\mathbb{Z}_{p}$
    $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$.
@@
    for every $c\in\overline{\mathbb{Q}_{p}}$ such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$, $v(s_{n+1}-c)\leq v(s_{n+1}-s_{n})$.
@@
    for every $c\in\overline{\mathbb{Q}_{p}}$ such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$, $v(s_{n}-c)<v(s_{n+1}-s_{n})$.
@@
Let $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ be a stacked sequence. Note that the sequence $\{v(s_{n+1}-s_{n})=\delta_{n}\}_{n\in\mathbb{N}}$ is strictly increasing, since $[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and $(s_{n},\delta_{n})$ is a minimal pair. In the original definition of distiguished sequence $E$ in [5], the sequence $\{\delta_{n}\}_{n\in\mathbb{N}}$ is unbounded, thus, in this case $E$ is a Cauchy sequence. In our setting we are not imposing that restriction; we show in Lemma 2.3 below that a stacked sequence is a pseudo-convergent sequence of transcendental type of $\overline{\mathbb{Q}_{p}}$.
The motivation for the terminology of these kind of sequences, is due to the following fact. For each $n\in\mathbb{N}$, for short we set $\Gamma_{n}=\Gamma_{s_{n}}$ and $k_{n}=k_{s_{n}}$ (i.e., the value group and the residue field of the valuation domain $O_{s_{n}}$ of $\mathbb{Q}_{p}(s_{n})$, respectively). By Remarks 1.6, $v(s_{n+1}-s_{n})>\delta_{K}(s_{n})$. Hence, by Theorem 1.3, we have $\Gamma_{n}\subseteq\Gamma_{n+1}$ and $k_{n}\subseteq k_{n+1}$. For each $n\in\mathbb{N}$, we set $e_{n}=e(\mathbb{Q}_{p}(s_{n})\mid\mathbb{Q}_{p})$ and $f_{n}=f(\mathbb{Q}_{p}(s_{n})\mid\mathbb{Q}_{p})$, the ramification index and the residue field degree of $O_{s_{n}}$ over $\mathbb{Z}_{p}$, respectively; we remark that $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=e_{n}f_{n}=d_{n}$ for each $n\in\mathbb{N}$, and since $\{d_{n}\}_{n\in\mathbb{N}}$ is unbounded by assumption, either $\{e_{n}\}_{n\in\mathbb{N}}$ is unbounded or $\{f_{n}\}_{n\in\mathbb{N}}$ is unbounded. Since $e_{n}\mid e_{n+1}$ for each $n\in\mathbb{N}$, $\{e_{n}\}_{n\in\mathbb{N}}$ is bounded if and only if $e_{n}=e$ for all $n\in\mathbb{N}$ sufficiently large. Similarly for $\{f_{n}\}_{n\in\mathbb{N}}$.
By Remarks 1.6, condition ii) is equivalent to $\delta_{n}=v(s_{n}-s_{n+1})=\delta_{K}(s_{n+1})$ (note that in general the inequality $\delta_{n}\leq\delta_{K}(s_{n+1})$ holds). In other words, among all the elements $c\in\overline{\mathbb{Q}_{p}}$ such that $[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]\leq[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$, $s_{n}$ is one of those which is closest to $s_{n+1}$.
@@
We prove now that $E$ is of transcendental type. Let $\alpha\in\overline{\mathbb{Q}_{p}}$. Then there exists $n\in\mathbb{N}$ such that $[\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Since $(s_{n},\delta_{n})$ is a minimal pair, $v(s_{n}-\alpha)<\delta_{n}$, so, in particular, $\alpha$ cannot be a pseudo-limit of $E$. This shows that $E$ has no pseudo-limits in $\overline{\mathbb{Q}_{p}}$, thus $E$ is of transcendental type. ∎
@@
    $[\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$.
@@
    the positive integer $[\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]-[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ is minimal.
Note that since $\mathbb{N}$ is well-ordered, condition iii) can be satisfied (that is, among the $\alpha\in\overline{\mathbb{Q}_{p}}$ satisfying i) and ii), we can find one which also satisfies iii) ). Since $\overline{v}_{E}(X-s_{n-1})=v(t_{m}-s_{n-1})$ for all $m$ sufficiently large, for such $m$’s we also have $\overline{v}_{E}(X-s_{n-1})<\overline{v}_{E}(X-t_{m})$. Moreover, without loss of generality, we may also assume that $[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ since $\{[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]\}_{m\in\mathbb{N}}$ is unbounded. This shows that the set $B_{n}$ is non-empty. Let $M_{E}(X,B_{n})=\{\overline{v}_{E}(X-\alpha)\mid\alpha\in B_{n}\}$, which is a subset of $\mathbb{Q}$. Let $\delta_{n}=\sup M_{E}(X,B_{n})$. Since each element of $B_{n}$ has the same degree over $\mathbb{Q}_{p}$, it follows that $B_{n}$ is contained in a finite extension $K$ of $\mathbb{Q}_{p}$. In particular, it follows as above that $M_{E}(X,B_{n})$ is upper bounded. Let $\delta_{n}=\sup M_{E}(X,B_{n})\in\mathbb{R}$. Next, we show that $M_{E}(X,B_{n})$ contains its upper bound (which is, in particular, a rational number). Suppose otherwise: then there exists a sequence $\{\alpha_{k}\}_{k\in\mathbb{N}}\subset B_{n}$ such that $\overline{v}_{E}(X-\alpha_{k})\nearrow\delta_{n}$. In particular, $\{\alpha_{k}\}_{k\in\mathbb{N}}$ would be a pseudo-convergent sequence of a finite extension of $\mathbb{Q}_{p}$ which is not Cauchy, which is impossible. Let $s_{n}\in B_{n}$ be such that $\overline{v}_{E}(X-s_{n})=\delta_{n}$. Note that $v_{p}(s_{n}-s_{n-1})=\overline{v}_{E}(s_{n}-X+X-s_{n-1})=\overline{v}_{E}(X-s_{n-1})=\delta_{n-1}$.
We now show that $(s_{n-1},s_{n})$ is distinguished. Clearly, $[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$.
Let $c\in\overline{\mathbb{Q}_{p}}$ be such that $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$.
If $[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$, then by the minimality of the degree of $s_{n}$, we have $\overline{v}_{E}(X-c)\leq\overline{v}_{E}(X-s_{n-1})=\delta_{n-1}$, so

### v3

@@ Introduction
Here, given a prime $p\in\mathbb{Z}$ and the DVR $\mathbb{Z}_{(p)}$ of $\mathbb{Q}$, we are interested in describing residually algebraic torsion extensions of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, that is, valuation domains $W$ of $\mathbb{Q}(X)$ lying above $\mathbb{Z}_{(p)}$ such that the residue field extension $W/M\supseteq\mathbb{Z}/p\mathbb{Z}$ is algebraic and the value group $\Gamma_{w}$ of the associated valuation $w$ to $W$ is contained in the divisible hull of the value group of $\mathbb{Z}_{(p)}$ (i.e., the rationals). These valuation domains arise naturally as overrings of rings of integer-valued polynomials and Dedekind domains between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ (see [9, 21]) and also in the description of closed subfields of $\mathbb{C}_{p}$ ([12]), the completion of an algebraic closure $\overline{\mathbb{Q}_{p}}$ of the field of $p$-adic numbers $\mathbb{Q}_{p}$. In the case $W$ is a DVR and the residue field extension is finite, then by [19, Theorem 2.5 & Proposition 2.2] there exists an element $\alpha$ in $\overline{\mathbb{Q}_{p}}$, transcendental over $\mathbb{Q}$, such that $W=\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\overline{\mathbb{Z}_{p}}\}$, where $\overline{\mathbb{Z}_{p}}$ is the absolute integral closure of $\mathbb{Z}_{p}$ (i.e., the integral closure of $\mathbb{Z}_{p}$ in $\overline{\mathbb{Q}_{p}}$; note that $\overline{\mathbb{Z}_{p}}$ is the valuation domain of the unique extension of $v_{p}$ to $\overline{\mathbb{Q}_{p}}$). In general, given a residually algebraic torsion extension $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$, there exists a pseudo-convergent sequence $E=\{s_{n}\}_{n\in\mathbb{N}}$ in $\overline{\mathbb{Q}_{p}}$ such that $W=\mathbb{Z}_{(p),E}=\{\phi\in\mathbb{Q}(X)\mid\phi(s_{n})\in\overline{\mathbb{Z}_{p}},\text{ for all sufficiently large }n\in\mathbb{N}\}$ (Proposition 2.24). One of the main result of this paper is to show that we may assume that $E$ is stacked (in a sense we make clear in §2; see Theorem 2.5). In particular, if $W$ is a DVR of $\mathbb{Q}(X)$ extending $\mathbb{Z}_{(p)}$ such that the extension of the residue fields is infinite algebraic, then there exists $\alpha$ in $\mathbb{C}_{p}\setminus\overline{\mathbb{Q}_{p}}$, such that $W=\mathbb{Z}_{(p),\alpha}=\{\phi\in\mathbb{Q}(X)\mid\phi(\alpha)\in\mathbb{O}_{p}\}$, where $\mathbb{O}_{p}$ is the completion of $\overline{\mathbb{Z}_{p}}$ (equivalently, $\mathbb{O}_{p}$ is the valuation domain of the unique extension of $v_{p}$ to $\mathbb{C}_{p}$; see Corollary 2.28). Necessarily, the (transcendental) extension $\mathbb{Q}_{p}(\alpha)/\mathbb{Q}_{p}$ has finite ramification.
@@
It is worth recalling that in [3, §5, 1., & Theorem 5.1] a residually algebraic torsion extension $W$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ is realized as the limit of a sequence of residually transcendental extensions $W_{n}$ of $\mathbb{Z}_{(p)}$ to $\mathbb{Q}(X)$ (i.e., the residue field extension of $W_{n}$ over $\mathbb{Z}_{(p)}$ is transcendental); moreover, for each $n\in\mathbb{N}$, $W_{n}$ is defined by a minimal pair $(s_{n},\delta_{n})$ (as explained in [3, p. 282]; for the definition of minimal pair see §1.2). Here, $W$ is realized as the valuation domain $\mathbb{Z}_{(p),E}$, where for each $n\in\mathbb{N}$, $(s_{n},\delta_{n}=v_{p}(s_{n+1}-s_{n}))$ is a minimal pair, too.
@@
As an application of the above results, we are able to complete the classification of the family of Dedekind domains $R$ between $\mathbb{Z}[X]$ and $\mathbb{Q}[X]$ started in [21]. In that paper we described the Dedekind domains of this family whose residue fields of prime characteristic are finite fields ([21, Theorem 2.17]); the description is obtained by means of the notion of rings of integer-valued polynomials over algebras. We also showed that, given a group $G$ which is the direct sum of a countable family of finitely generated abelian groups, there exists a Dedekind domain $R$ with finite residue fields of prime characteristic, $\mathbb{Z}[X]\subset R\subseteq\mathbb{Q}[X]$, with class group $G$ ([21, Theorem 3.1]).
@@ 1.2 Distinguished pairs
By Krasner’s Lemma, for each of the relevant $c$’s we have $v(a-c)\leq\omega_{K}(a)$. Note that the ramification index of $K(a,c)$ over $K$ is (strictly) bounded by $K(a):K]^{2}$. In particular, since the value v(ac)v(a-c) belongs to $\Gamma_{a-c}$, it follows that there exists $N\in\mathbb{N}$, independent from each of the above $c$’s, such that $Nv(a-c)\in\Gamma_{v}\cong\mathbb{Z}$. Hence, the set $\{v(a-c)\mid c\in\overline{K},[K(c):K]<[K(a):K]\}$ (which is a subset of $\Gamma_{\overline{v}}$) is bounded from above and its elements have bounded torsion. It follows that this set has a maximum, which is equal to $\delta_{K}(a)$ by its very definition. ∎
@@
Let $(a,\delta)\in\overline{K}\times\Gamma_{\overline{v}}$. We say that $(a,\delta)$ is a minimal pair if for every $c\in\overline{K}$ such that $[K(c):K]<[K(a):K]$ we have $v(a-c)<\delta$.
In other words, $(a,\delta)$ is a minimal pair if for every $b\in B(a,\delta)=\{x\in\overline{K}\mid v(a-x)\geq\delta\}$, we have $K(b):K]\geq[K(a):K]$ (i.e., $a$ is a ’center’ of the ball $B(a,\delta)$ of minimal degree). By Lemma 1.2, $(a,\delta)$ is a minimal pair if and only if $\delta>\delta_{K}(a)$. In particular, if $\delta>\omega_{K}(a)$, then $(a,\delta)$ is a minimal pair.
@@
Let $(b,a)$ be a distinguished pair. Note that i) and ii) imply that $v(a-b)=\delta_{K}(a)$. In fact, by i) and ii), it immediately follows that the inequality $\leq$ holds. Conversely, by ii) we also have that $v(a-b)\geq v(a-c)$ for all $c$ such that $[K(c):K]<[K(a):K]$, that is, $v(a-b)\geq\delta_{K}(a)$.
@@
    for all $c\in\overline{K}$ such that $[K(c):K]<[K(b):K]$ then v(bc)<v(ab)v(b-c)<v(a-b).
which precisely says that (b,v(ab))(b,v(a-b)) is a minimal pair with respect to $K$. In fact, if iii) holds and $c\in\overline{K}$ is such that $[K(c):K]<[K(b):K]$ then v(bc)=v(ba+ac)=v(ac)<v(ab)v(b-c)=v(b-a+a-c)=v(a-c)<v(a-b). Similarly, iii’) implies iii). Note also that iii’) is equivalent to
@@
In particular, by the above Theorem, $\Gamma_{b}\subseteq\Gamma_{a}$, $k_{b}\subseteq k_{a}$ and $K(b):K]\mid[K(a):K]$.
@@ 2 Stacked pseudo-convergent sequences of $\overline{\mathbb{Q}_{p}}$
Let $\mathbb{P}\subset\mathbb{Z}$ be the set of prime numbers and let $p\in\mathbb{P}$ be a fixed prime. We let $\mathbb{Z}_{(p)}$ be the localization of $\mathbb{Z}$ at the prime ideal $\mathbb{Z}, p\mathbb{Z}_{p}$ the ring of $p$-adic integers and $\mathbb{Q}_{p}$ its field of fractions, the field of $p$-adic numbers. If $v_{p}$ denotes the usual $p$-adic valuation, then $\mathbb{Z}_{p}$ ($\mathbb{Q}_{p}$, respectively) is the completion of $\mathbb{Z}$ ($\mathbb{Q}$, respectively) with respect to the $p$-adic valuation. We denote by $\overline{\mathbb{Q}_{p}}$ a fixed algebraic closure of $\mathbb{Q}_{p}$ and still denote the unique extension of $v_{p}$ to $\overline{\mathbb{Q}_{p}}$ by $v_{p}$. Note that $\overline{\mathbb{Q}_{p}}$ is a rank one non-discrete valued field with valuation domain denoted by $\overline{\mathbb{Z}_{p}}$, the integral closure of $\mathbb{Z}_{p}$ in $\overline{\mathbb{Q}_{p}}$. We will use the well-known fact that $\mathbb{Q}_{p}$ has only finitely many extensions of a given degree (see for example [17, Corollary 2, Chapter V, p. 202]).
@@ 2.1 Residually algebraic torsion extensions of $\mathbb{Z}_{p}$
    $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$.
@@
    for every $c\in\overline{\mathbb{Q}_{p}}$ such that $\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$, $v(s_{n+1}-c)\leq v(s_{n+1}-s_{n})$.
@@
    for every $c\in\overline{\mathbb{Q}_{p}}$ such that $\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$, $v(s_{n}-c)<v(s_{n+1}-s_{n})$.
@@
Let $E=\{s_{n}\}_{n\in\mathbb{N}}\subset\overline{\mathbb{Q}_{p}}$ be a stacked sequence. Note that the sequence $\{v(s_{n+1}-s_{n})=\delta_{n}\}_{n\in\mathbb{N}}$ is strictly increasing, since $\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$ and $(s_{n},\delta_{n})$ is a minimal pair. In the original definition of distiguished sequence $E$ in [5], the sequence $\{\delta_{n}\}_{n\in\mathbb{N}}$ is unbounded, thus, in this case $E$ is a Cauchy sequence. In our setting we are not imposing that restriction; we show in Lemma 2.3 below that a stacked sequence is a pseudo-convergent sequence of transcendental type of $\overline{\mathbb{Q}_{p}}$.
The motivation for the terminology of these kind of sequences, is due to the following fact. For each $n\in\mathbb{N}$, for short we set $\Gamma_{n}=\Gamma_{s_{n}}$ and $k_{n}=k_{s_{n}}$ (i.e., the value group and the residue field of the valuation domain $O_{s_{n}}$ of $\mathbb{Q}_{p}(s_{n})$, respectively). By Remarks 1.6, $v(s_{n+1}-s_{n})>\delta(s_{n})$. Hence, by Theorem 1.3, we have $\Gamma_{n}\subseteq\Gamma_{n+1}$ and $k_{n}\subseteq k_{n+1}$. For each $n\in\mathbb{N}$, we set $e_{n}=e(\mathbb{Q}_{p}(s_{n})\mid\mathbb{Q}_{p})$ and $f_{n}=f(\mathbb{Q}_{p}(s_{n})\mid\mathbb{Q}_{p})$, the ramification index and the residue field degree of $O_{s_{n}}$ over $\mathbb{Z}_{p}$, respectively; we remark that $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]=e_{n}f_{n}=d_{n}$ for each $n\in\mathbb{N}$, and since $\{d_{n}\}_{n\in\mathbb{N}}$ is unbounded by assumption, either $\{e_{n}\}_{n\in\mathbb{N}}$ is unbounded or $\{f_{n}\}_{n\in\mathbb{N}}$ is unbounded. Since $e_{n}\mid e_{n+1}$ for each $n\in\mathbb{N}$, $\{e_{n}\}_{n\in\mathbb{N}}$ is bounded if and only if $e_{n}=e$ for all $n\in\mathbb{N}$ sufficiently large. Similarly for $\{f_{n}\}_{n\in\mathbb{N}}$.
By Remarks 1.6, condition ii) is equivalent to $\delta_{n}=v(s_{n}-s_{n+1})=\delta(s_{n+1})$ (note that in general the inequality $\delta_{n}\leq\delta(s_{n+1})$ holds). In other words, among all the elements $c\in\overline{\mathbb{Q}_{p}}$ such that $\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]\leq[\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n+1}):\mathbb{Q}_{p}]$, $s_{n}$ is one of those which is closest to $s_{n+1}$.
@@
We prove now that $E$ is of transcendental type. Let $\alpha\in\overline{\mathbb{Q}_{p}}$. Then there exists $n\in\mathbb{N}$ such that $\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$. Since $(s_{n},\delta_{n})$ is a minimal pair, $v(s_{n}-\alpha)<\delta_{n}$, so, in particular, $\alpha$ cannot be a pseudo-limit of $E$. This shows that $E$ has no pseudo-limits in $\overline{\mathbb{Q}_{p}}$, thus $E$ is of transcendental type. ∎
@@
    $\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$.
@@
    the positive integer $\mathbb{Q}_{p}(\alpha):\mathbb{Q}_{p}]-[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ is minimal.
Note that since $\mathbb{N}$ is well-ordered, condition iii) can be satisfied (that is, among the $\alpha\in\overline{\mathbb{Q}_{p}}$ satisfying i) and ii), we can find one which also satisfies iii) ). Since $\overline{v}_{E}(X-s_{n-1})=v(t_{m}-s_{n-1})$ for all $m$ sufficiently large, for such $m$’s we also have $\overline{v}_{E}(X-s_{n-1})<\overline{v}_{E}(X-t_{m})$. Moreover, without loss of generality, we may also assume that $\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$ since $\{[\mathbb{Q}_{p}(t_{m}):\mathbb{Q}_{p}]\}_{m\in\mathbb{N}}$ is unbounded. This shows that the set $B_{n}$ is non-empty. Let $M_{E}(X,B_{n})=\{\overline{v}_{E}(X-\alpha)\mid\alpha\in B_{n}\}$, which is a subset of $\mathbb{Q}$. Let $\delta_{n}=\sup M_{E}(X,B_{n})$. Since each element of $B_{n}$ has the same degree over $\mathbb{Q}_{p}$, it follows that $B_{n}$ is contained in a finite extension $K$ of $\mathbb{Q}_{p}$. In particular, it follows as above that $M_{E}(X,B_{n})$ is upper bounded. Let $\delta_{n}=\sup M_{E}(X,B_{n})\in\mathbb{R}$. Next, we show that $M_{E}(X,B_{n})$ contains its upper bound (which is, in particular, a rational number). Suppose otherwise: then there exists a sequence $\{\alpha_{k}\}_{k\in\mathbb{N}}\subset B_{n}$ such that $\overline{v}_{E}(X-\alpha_{k})\nearrow\delta_{n}$. In particular, $\{\alpha_{k}\}_{k\in\mathbb{N}}$ would be a pseudo-convergent sequence of a finite extension of $\mathbb{Q}_{p}$ which is not Cauchy, which is impossible. Let $s_{n}\in B_{n}$ be such that $\overline{v}_{E}(X-s_{n})=\delta_{n}$. Note that $v_{p}(s_{n}-s_{n-1})=\overline{v}_{E}(s_{n}-X+X-s_{n-1})=\overline{v}_{E}(X-s_{n-1})=\delta_{n-1}$.
We now show that $(s_{n-1},s_{n})$ is distinguished. Clearly, $\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$.
Let $c\in\overline{\mathbb{Q}_{p}}$ be such that $\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]<[\mathbb{Q}_{p}(s_{n}):\mathbb{Q}_{p}]$.
If $\mathbb{Q}_{p}(c):\mathbb{Q}_{p}]>[\mathbb{Q}_{p}(s_{n-1}):\mathbb{Q}_{p}]$, then by the minimality of the degree of $s_{n}$, we have $\overline{v}_{E}(X-c)\leq\overline{v}_{E}(X-s_{n-1})=\delta_{n-1}$, so